please DO NOT use mesh analysis for it I need Z11 .. Z22 and just use the formula , thanks
Problem 2 Determine the z parameters of the circuits below. 202 WWW O 102 AMM 322 49 WW
Answer #1

Solution I w t. I 3л 늘 Am 40 → Z parameters V = Z, F1 + Z2I2 V2 = Z21. It R22 Iz » Z = Vi 212= V T 11,-0 I, /11=0 Za Z22= V₂☆ for 212 4. Zzz & Io. Iz 2 1,=0 Iz + Ty + + = (1+2+4), 230 V, 31 yu Ir + 3 + Iy = (I) = 3 Ir 2 10 7+3 =2158.1 Jy = .0.37 10د Now 03 = درد وہ: 0,332 = 7. = در 1 a = = = = .2 = 5/3 ر2 - زده دار 1 0.3 ۰۹ X) 32 0.3 اد2 اد ر - Thank your CS scanned with